OpenStudy (anonymous):

The surface area of the cylinder is 219in2 and the height of the cylinder is 12 inches. What is the diameter of the cylinder?

OpenStudy (anonymous):

\[SA=2\pi rh+2\pi r^2\]

OpenStudy (anonymous):

That could help (the formula)

OpenStudy (anonymous):

izit 2.90?

OpenStudy (anonymous):

aleroth are u almost done

OpenStudy (anonymous):

219=2πrh+2π(r^2) 219=2πr(12)+2π(r^2) 219=24πr+2π(r^2) 219=2π(r^2)+24πr 219=2π ((r^2)+12r) 219+72π=2π ((r^2)+12r +36) (219+72π)/2π = (r+6)^2 squareroot((219+72π)/2π) = r+6 r= 2.417 approximately

OpenStudy (anonymous):

SA = 2*pi*r^2 + h(2*pi*r) = 219 2*pi*r^2 + 12*2*pi*r* = 219 2*pi*r^2 + 24*pi*r = 219 pi*r^2 + 12*pi*r = 109.5 r^2 + 12r = 109.5/pi If pi = 3.14 r^2 + 12r = 34.8726 r^2 + 12r - 34.8726 = 0 r=-6+-sqrt(6^2+34.8726) r=-6+-sqrt(70,8726) r=-6+-8.418586579705646 r=2.4185865797056465 d=4.837173159411293